# SimMerge: Learning to Select Merge Operators from Similarity Signals

### v1

@@ 4.2 From Pairwise Selection to Multi-Way Merges
We next ask whether a selector trained only on 2-way merges can be lifted to the multiway setting. We encode a $k$-way merge plan $\pi$ as an order-aware sequence of pairwise feature blocks $X(\pi,t)$ and reuse the same similarity encoder and selector architectures to score candidate plans, without any additional $k$-way supervision.
Figure 4: Overall relative performance of 2-, 3-, and 4-way merges, reported as percentage change vs. the task expert (top) and vs. auxiliary models (bottom), macro-averaged over all tasks. SimMerge consistently improves over auxiliaries while limiting degradation relative to experts as the number of merged models increases.
Figure 4 summarizes the effect of increasing the number of merged models from $k=2$ to $k=3$ and $k=4$, macro-averaged over all tasks. The left panel shows percentage change in performance relative to the task expert, and the right panel shows the corresponding change relative to the auxiliary models. As expected, increasing $k$ makes merging more challenging for all methods: degradation relative to the expert grows and gains over auxiliaries shrink. Across all values of $k$, however, SimMerge achieves the best trade-off on both axes. For instance, it stays within $14.5\%$ to $17.9\%$ of expert performance as $k$ increases from 2 to 4, while still improving over auxiliaries by $28.0\%$, $17.7\%$, and $15.7\%$ at $k=2,3,4$. These results mirror the pairwise trends in Section 4.1 and suggest that similarity-driven selection learned from 2-way merges transfers to multiway merge plans. SimMerge remains the best-performing method on both axes across all values of $k$. Detailed per-task results for 3- and 4-way merges are provided in Appendix F.2, where we observe the same qualitative behavior across code, math, multilingual, and RAG.
@@ 4.3 Effect of Merge Order
For 3-way merges, the merge operator is non-associative, so the order in which models are merged can change the final performance. We therefore compare the merge order chosen by our learned selector to a random-order baseline. To isolate the effect of ordering from operator choice, we keep the same sequence of merge operators selected by SimMerge fixed and only randomize the permutation of models. This isolates the effect of the merge order itself, independent of the choice of operators.
Figure 5: Performance of three-way merges across Code, Math, Multilingual, and RAG tasks, measured as the percentage of the expert–auxiliary performance gap that is closed. Auxiliary performance corresponds to 0%, expert performance to 100%.
Figure 5 compares SimMerge’s selected order to a random order baseline using the same operator sequence, reported in terms of GapClosed. The numeric labels denote the improvement of the learned ordering over the random baseline. Optimized ordering delivers consistent gains across tasks: SimMerge closes a $47\%$ gap on Code, $16\%$ on Math, $11\%$ on Multilingual, and $21\%$ on RAG compared to a random ordering. These results indicate that similarity signals are useful not only for selecting merge operators, but also for selecting the merge order, with especially large order effects in Code and RAG.
@@ 4.4 Scaling to a  [^1] B-Parameter Model
Figure 6: Overall percentage change in performance for 3-way merges on a  [^1] B-parameter model, relative to the task expert (left) and to the auxiliary models (right), macro-averaged over all tasks. Each bar corresponds to using one operator uniformly for all merges; SimMerge (green) selects among Linear, SLERP, and TIES using similarity features and yields the smallest degradation relative to the expert and the largest gains over auxiliary models, despite being trained only on 7B-parameter checkpoints.
Figure 6 reports the overall percentage change in performance relative to the task expert (left panel) and to the auxiliary baseline (right panel), averaged over tasks. On the  [^1] B model, SimMerge again achieves the best trade-off between expert and auxiliary performance: it reduces degradation relative to the expert to 7.8-7.8% compared to 14.5-14.5% (Linear), 12.4-12.4% (SLERP), and 13.9-13.9% (TIES). At the same time, it achieves the largest improvement over auxiliary models at +42.7+42.7% (SimMerge) versus +34.8+34.8% (Linear), +36.0+36.0% (SLERP), and +33.4+33.4% (TIES). Appendix F.5 provides a domain-level breakdown for 3-way merges at  [^1] B using both $\Delta_{\text{expert}}/\Delta_{\text{aux}}$ and GapClosed, and shows the same qualitative pattern across Code, Math, Multilingual, and RAG.
Taken together with the pairwise and multiway results in Sections 4.1 and 4.2, these findings demonstrate that predictive merge selection transfers from 7B to  [^1] B parameters without retraining while providing consistent gains over strong fixed merge operators.
@@ 4.5 Online Learning
Figure 7: Cumulative regret vs. an oracle policy of 3-way merges. The uniform random baseline accumulates regret nearly linearly. Both neural-linear bandits, LinUCB and LinTS, learn to select better operators, with LinTS achieving consistently lower regret.
Figure 7 shows cumulative regret relative to the oracle over 60 rounds. Uniform random accumulates regret roughly linearly. Both contextual bandits learn quickly and substantially reduce regret, but LinTS dominates LinUCB across the entire horizon: it converges faster and tracks the oracle more closely, indicating that the similarity features are informative enough to support efficient exploration and exploitation for operator selection.
Figure 8: Percentage change in performance for 3-way merges under three online policies, relative to the task expert (left) and to auxiliary models (right), across code, math, multilingual, RAG, instruct, and the overall average. The oracle chooses the best operator per round. LinTS (our bandit SimMerge) closely matches the oracle and consistently outperforms the uniform random baseline, achieving smaller degradation vs. experts and larger gains vs. auxiliaries.
To connect regret to downstream quality, Figure 8 reports the final percentage change in performance for three representative policies: uniform random, LinTS, and oracle, relative to the task expert (left) and to the auxiliary baseline (right), aggregated over 3-way merges on Code, Math, Multilingual, RAG, Instruct, and the overall macro-average. LinTS closely tracks the oracle on both axes and consistently improves over the uniform random baseline. For example, on the overall macro-average it reaches $18.8\%$ relative to the expert while improving by $13.7\%$ over the auxiliary baseline. These results show that the same similarity features can support a practical online policy via Thompson sampling, enabling adaptation to distribution shift while retaining much of the benefit of full-information operator choice.
@@ 5 Related Work
Automating merge hyperparameters. Closer to our goal are methods that tune merge hyperparameters rather than the base operator itself. Fisher merging [Matena & Raffel, 2022] uses data-derived Fisher information to compute a weighted average. AdaMerge [Yang et al., 2024] chooses task-vector coefficients to minimize merge-model entropy on unlabeled data. Akiba et al. [2025] apply evolutionary search to optimize merge recipes, iteratively refining coefficients and layer-wise schedules. These approaches still operate within a fixed operator family and typically require multiple merge-and-evaluate steps per configuration.
@@ Appendix D Propagation of Similarity Metrics to Multi-Way Plans
$(a{+}b,\,c):\quad G_{L}=(1-\alpha)P_{a}+\alpha P_{b},\;G_{R}=P_{c},\qquad\text{and}\qquad(c,\,a{+}b):\quad G_{L}=P_{c},\;G_{R}=(1-\alpha)Q_{a}+\alpha Q_{b},$
@@ F.3 Per-Task Trends Across Merge Sizes
Auxiliary-relative trends (Math and RAG). Figure 11 plots $\Delta_{\text{aux}}$ for Math and RAG. As $k$ grows, auxiliary gains can shrink or even become negative for fixed operators, indicating that naive multi-way merges can underperform the auxiliary baseline. This behavior is especially visible in RAG, where interference is strong and fixed operators often yield weak or negative auxiliary improvements. In contrast, SimMerge more reliably maintains positive (or near-zero) auxiliary gains across merge sizes, suggesting better retention of off-domain capability while still limiting expert degradation (Figure 10).

### v2

@@ 4.2 From Pairwise Selection to Multi-Way Merges
Figure 4: Overall relative performance of 2-, 3-, and 4-way merges, reported as percentage change vs. the task expert (left) and vs. auxiliary models (right), macro-averaged over all tasks. SimMerge consistently improves over auxiliaries while limiting degradation relative to experts as the number of merged models increases.
We next ask whether a selector trained only on 2-way merges can be lifted to the multiway setting. We encode a $k$-way merge plan $\pi$ as an order-aware sequence of pairwise merges and reuse the same similarity encoder and selector architectures to score candidate plans, without any additional $k$-way supervision.
Figure 4 summarizes the effect of increasing the number of merged models from $k=2$ to $3$ and $4$, macro-averaged over all tasks. The left panel shows percentage change in performance relative to the task expert, and the right panel shows the corresponding change relative to the auxiliary models. As expected, increasing $k$ (merging more models by adding more auxiliaries) makes merging more challenging for all methods: degradation relative to the expert grows and gains over auxiliaries shrink. Across all values of $k$, however, SimMerge consistently produces the strongest merges among the compared methods, preserving more expert performance while achieving larger improvements over auxiliaries. For instance, its macro-averaged change relative to the expert is $14.5\%$, $16.0\%$, and $17.9\%$ for $k=2,3,4$, while the change relative to auxiliaries remains positive at $28.0\%$, $17.7\%$, and $15.7\%$, respectively. These results mirror the pairwise trends in Section 4.1 and suggest that similarity-driven selection learned from 2-way merges transfers to multiway merge plans.
Detailed per-task results for 3- and 4-way merges are provided in Appendix F.2, where we observe the same behavior across code, math, multilingual, and RAG.
@@ 4.3 Effect of Merge Order
For 3-way and 4-way merges, the merge operation is non-associative, so the order in which models are merged can change the final performance. We therefore compare the merge order chosen by our learned selector to a random-order baseline. To isolate the effect of ordering from operator choice, we keep the same sequence of merge operators selected by SimMerge fixed and only randomize the merge order.
Figure 5: Performance of 3-way merges across Code, Math, Multilingual, and RAG tasks, measured as the percentage of the expert–auxiliary performance gap that is closed. Auxiliary performance corresponds to 0%, expert performance to 100%.
Figure 5 compares SimMerge’s selected order to a random-order baseline using the same operator sequence, reported in terms of GapClosed. The numeric labels denote the improvement (percentage points) of the learned ordering over the random baseline. Optimized ordering of SimMerge delivers consistent gains across tasks, improving GapClosed by +47+47 points on Code, +16+16 on Math, +11+11 on Multilingual, and +21+21 on RAG compared to a random ordering. These results indicate that similarity signals are useful not only for selecting merge operators, but also for selecting merge order, with especially large order effects in Code and RAG.
@@ 4.4 Scaling to a  [^1] B-Parameter Model
Figure 6: 3-way merging at  [^1] B: macro-averaged performance change vs. the task expert (left) and auxiliary models (right). SimMerge yields the smallest expert degradation and largest gains over auxiliaries, despite training only on 7B checkpoints.
Figure 6 reports the overall percentage change in performance relative to the task expert (left panel) and to the auxiliary baseline (right panel), averaged over tasks. On the  [^1] B model, SimMerge again achieves the best trade-off between expert and auxiliary performance: it reduces degradation relative to the expert to 7.8-7.8% compared to 14.5-14.5% (Linear), 12.4-12.4% (Slerp), and 13.9-13.9% (TIES). At the same time, it achieves the largest improvement over auxiliary models at +42.7+42.7% (SimMerge) versus +34.8+34.8% (Linear), +36.0+36.0% (Slerp), and +33.4+33.4% (TIES). Appendix F.5 provides a domain-level breakdown for 3-way merges at  [^1] B using both $\Delta_{\text{expert}}/\Delta_{\text{aux}}$ and GapClosed, and shows the same qualitative pattern across Code, Math, Multilingual, and RAG. Taken together with the pairwise and multiway results in Sections 4.1 and 4.2, these findings demonstrate that predictive merge selection transfers from 7B to  [^1] B parameters without retraining while providing consistent gains over strong fixed merge operators.
@@ 4.5 Online Learning
Figure 7: Cumulative regret relative to the oracle over 60 rounds for 3-way merges. Uniform random accumulates regret roughly linearly; both contextual bandits reduce regret quickly, with LinTS consistently lower than LinUCB.
Figure 8 (left) shows cumulative regret relative to the oracle over 60 rounds. Uniform random accumulates regret roughly linearly. Both contextual bandits learn quickly and substantially reduce regret, but LinTS dominates LinUCB across the entire horizon: it converges faster and tracks the oracle more closely, indicating that the similarity features are informative enough to support efficient exploration and exploitation for operator selection.
Figure 8: Final percentage change in performance for 3-way merges under uniform random, LinTS, and oracle policies, relative to the task expert (left axis) and auxiliary baseline (right axis), aggregated across domains and macro-average.
To connect regret to downstream quality, Figure 8 (right) reports the final percentage change in performance for three representative policies – uniform random, LinTS, and oracle – relative to the task expert and to the auxiliary baseline, aggregated over 3-way merges on Code, Math, Multilingual, RAG, Instruct, and the overall macro-average. LinTS closely tracks the oracle on both axes and consistently improves over uniform random. For example, on the macro-average it reaches $18.8\%$ relative to the expert while improving by $13.7\%$ over the auxiliary baseline. These results show that the same similarity features can support a practical online policy via Thompson sampling, enabling adaptation to distribution shift while retaining much of the benefit of full-information operator choice.
@@ 5 Related Work
Automating merge hyperparameters. Closer to our goal are methods that tune merge hyperparameters rather than the base operator itself. Fisher merging [Matena & Raffel, 2022] uses data-derived Fisher information to compute a weighted average. AdaMerge [Yang et al., 2024] chooses task-vector coefficients to minimize merge-model entropy on unlabeled data. Khalifa et al. [2024] optimize linear merging coefficients over large pools of checkpoints by recycling checkpoints from previous runs to mitigate task tradeoffs, but still rely on repeated merge-and-evaluate iterations. Su & Geiping [2025] propose an automated model merging framework that uses multi-fidelity optimization to search over merge configurations, and introduce fine-grained search spaces such as layer-wise fusion and depth-wise integration to reduce search cost while improving single- and multi-objective performance. Akiba et al. [2025] similarly apply evolutionary search to optimize merge recipes, iteratively refining coefficients and layer-wise schedules. These approaches still operate within a fixed operator family and typically require multiple merge-and-evaluate steps per configuration.
@@ Appendix D Propagation of Similarity Metrics to Multi-Way Plans
$(a{+}b,\,c):\quad G_{L}=(1-\alpha)P_{a}+\alpha P_{b},\;G_{R}=P_{c},\quad\text{and}\quad(c,\,a{+}b):\quad G_{L}=P_{c},\;G_{R}=(1-\alpha)Q_{a}+\alpha Q_{b},$
@@ F.3 Per-Task Trends Across Merge Sizes
Auxiliary-relative trends (Math and RAG). Figure 11 plots $\Delta_{\text{aux}}$ for Math and RAG. As $k$ grows, auxiliary gains can shrink or even become negative for fixed operators, indicating that naive multi-way merges can underperform the auxiliary baseline. This behavior is especially visible in RAG, where interference is strong and fixed operators often yield weakly positive or negative auxiliary percentage change. In contrast, SimMerge more reliably maintains positive (or near-zero) auxiliary gains across merge sizes, suggesting better retention of off-domain capability while still limiting expert degradation (Figure 10).